A dolgozattal vizsgajogot nem szerzettek névsora

Breszkócs Péter Lajos, Husiová Anikó, Kiss Annamária, Kovács György, Kovács Szilvia, Kovács Zoltán, Kovácsné, Szilágyi Annamária, Palcsesz Viktória, Péter Antal, Tóth István Márk.